OpenStudy (anonymous):

At a rock concert, the sound intensity 1.0 m in front of the bank of loudspeakers is 0.10 W/m^2 . A fan is 30 m from the loudspeakers. Her eardrums have a diameter of 8.4 mm. How much energy is transferred to each eardrum in one second?

OpenStudy (anonymous):

Intensity decreases as 1/d^2 from source. thus at the position of ear it would be: \[0.1W/900m^2\] multiply this with the area of the ear \[\pi r^2=3.14*(0.0084m)^2\] you will get power of sound , which is energy per second \[(0.1W/900m^2)*3.14*(0.0084m)^2=2.46*10^{-8}W\]